OpenStudy (anonymous):

Chain link fence is used to make an animal pen of one rectangular section. There is 220 feet of chain link fence. The enclosed area is 3000 square feet. One side is longer than the other. Find the dimensions.

OpenStudy (anonymous):

A rectangular area can have a length and width designated. Let L = length and W = width. From the perimeter, we know that: Perimeter = P = 220ft = 2W + 2L P/2= 110ft = W + L The formula for a rectangle's area tells us that: Area = A = 3000ft^2 = W * L This can be solved by substituting: W = 110 - L 3000ft^2 = (110 - L) * L L^2 - 110L + 3000 = 0 (L - 60)(L-50) = 0 L = 50ft or L = 60ft Therefore, length and width are 50ft and 60ft, one way or the other. As length and width are arbitrarily defined in this process, solving the quadratic yields both side lengths.